Maggie’s brother is 4 years younger than three times her age . The sum of their age is 16 . How old is Maggie ?
Answer by ikleyn(52800)

.

Let x be the Maggie's age.


Then the brother's age is  3x - 4.


The total is


    x + (3x-4) = 16.


Simplify and solve for x.


    4x - 4 = 16

    4x     = 16 + 4

    4x     = 20

     x     = 20/4 = 5.


ANSWER.  Maggie is 5 years old.

Solved.